OpenStudy (anonymous):

Without using a calculator, find the exact value of cos^-1(cos(17pi/5)). Thanks for the help :)

OpenStudy (whpalmer4):

\[\cos^{-1}[\cos{(\frac{17\pi}{5})]}=\] Hopefully you recognize that you have what you started with, but reduced to be within the normal range. How many times can you subtract 2pi from 17pi/5 and stay > 0?

OpenStudy (anonymous):

your job is only to find the angle in the interval \([0,\pi]\) that has the same cosine as \(\frac{17\pi}{5}\)

OpenStudy (anonymous):

@whpalmer4 i think this one is a little trickier than that, because if you subtract \(2\pi\) you get \(\frac{7\pi}{5}\) which is still not quite good enough

OpenStudy (whpalmer4):

Wasn't intending that to be the whole solution, just a starting point

OpenStudy (whpalmer4):

And perhaps not artfully described, at that!

OpenStudy (anonymous):



OpenStudy (anonymous):

speaking of "art" that is truly ugly

OpenStudy (whpalmer4):

I'm surprised you don't have a library of drawings to attach by this point!

OpenStudy (anonymous):

actually i usually attach this but it is not helpful here

OpenStudy (anonymous):

find \(\frac{17\pi}{5}\) in my ugly picture it is this 

OpenStudy (anonymous):

now move directly north, so the cosine will be the same, but will be in the interval \([0,\pi]\)

OpenStudy (anonymous):

2pi/3